# Minimal Torsion Curves in Geometric Isogeny Classes

### v1

@@ 3.3. Image of Galois Representations Under Isogeny
    If $\operatorname{im}\rho_{E_{1},\ell^{\infty}}=\pi^{-1}(\operatorname{im}\rho_{E_{1},\ell^{k_{1}}})$ for $k_{1}\in\mathbb{Z}^{+}$, then $\operatorname{im}\rho_{E_{2},\ell^{\infty}}=\pi^{-1}(\operatorname{im}\rho_{E_{2},\ell^{k_{1}+r}})$. Here, $\pi$ denotes the reduction map from $\operatorname{GL}_{2}(\mathbb{Z}_{\ell})$ to $\operatorname{GL}_{2}(\mathbb{Z}/\ell^{k_{1}}\mathbb{Z})$ or $\operatorname{GL}_{2}(\mathbb{Z}/\ell^{k_{1}+r\mathbb{Z}})$, respectively.
@@
If we can show that $M\in\operatorname{im}\rho_{E_{2},\ell^{k_{1}+r+1}}$, then $\operatorname{im}\rho_{E_{2},\ell^{\infty}}=\pi^{-1}(\operatorname{im}\rho_{E_{2},\ell^{k_{1}+r}})$ by, e.g., [6, Proposition 3.5]. By the first paragraph, we may assume
@@
Suppose there exists an isogeny $\varphi:E\rightarrow E^{\prime}$ defined over $\overline{\mathbb{Q}}$ which is cyclic of order $\ell^{r}$. Thus there exists a point $P\in E$ of order $\ell^{r}$ such that $E^{\prime}\cong E/\langle P\rangle$. Let $F\coloneqq\mathbb{Q}(\langle P\rangle)$, and let $\psi:E\rightarrow E/\langle P\rangle$ be the induced $F$-rational isogeny. By Proposition 3.5, the $\ell$-adic Galois representation of $E/\langle P\rangle$ — and hence, the degrees of all closed points on $X_{1}(\ell^{k})$ associated to $E/\langle P\rangle$ — is determined by $\operatorname{im}\rho_{E/F,\ell^{\infty}}$. Since $\operatorname{im}\rho_{E/F,\ell^{\infty}}$ can be obtained from $\operatorname{im}\rho_{E/\mathbb{Q},\ell^{\infty}}$ and the definition of $F$, the result follows. ∎
@@ 4. Properties of Minimal Torsion Curves
Let $\mathcal{E}$ be a rational $\overline{\mathbb{Q}}$-isogeny class of elliptic curves. In this section, we show that for a fixed positive integer $N$, there are finitely many minimal torsion curves in $\mathcal{E}$ for $X_{1}(N)$. In some cases, there is a unique minimal torsion curve, but this should not be expected in general (see Remark 4.2). Next, we suppose $\mathcal{E}$ is non-CM, and let $E_{0}\in\mathcal{E}$ be an elliptic curve with rational $j$-invariant. Fix $\ell$ prime. In Proposition 4.3 we show that for any $k\in\mathbb{Z}^{+}$, there exists a minimal torsion curve $E\in\mathcal{E}$ for $X_{1}(\ell^{k})$ such that the degree of the isogeny from $E$ to $E_{0}$ is at most $C$, where $C$ is a constant that does not depend on $k$. This implies Proposition 4.3, as stated in the introduction.
@@ 4.1. Minimal Torsion Curves for Fixed Modular Curve
Let $d$ be the minimal degree of a point on $X_{1}(N)$ associated to $\mathcal{E}$, and let $j_{\text{min}}$ be the $j$-invariant of a minimal torsion curve for $\mathcal{E}$. Then $[\mathbb{Q}(j_{\text{min}}):\mathbb{Q}]\leq d.$ If $\mathcal{E}$ is CM, then the conclusion follows as there are only finitely many CM $j$-invariants in an extension of bounded degree (since there are only finitely many imaginary quadratic fields—and hence imaginary quadratic orders—of a given class number [26], this is a consequence of [13, Theorem 11.1 & Proposition 13.2]). So suppose $\mathcal{E}$ is non-CM.
Let $E_{\text{min}}\in\mathcal{E}$ be an elliptic curve with $j(E_{\text{min}})=j_{\text{min}}$. Since $\mathcal{E}$ is rational, there exists an elliptic curve $E_{0}/\mathbb{Q}\in\mathcal{E}$ with an isogeny $\varphi:E_{0}\rightarrow E_{\text{min}}$ defined over $\overline{\mathbb{Q}}$. We may assume $\varphi$ is cyclic of degree $n$ by [14, Lemma A.1]. By [11, Proposition 3.2] or [14, Corollary A.5], the field of moduli of this isogeny is
$\mathbb{Q}(j_{\text{min}},j(E_{0}))=\mathbb{Q}(j_{\text{min}}).$
Since isogenies are twist invariant, it follows that $E_{0}$ attains a rational cyclic $n$-isogeny over $\mathbb{Q}(j_{\text{min}})$, a number field of degree at most $d$.
By Serre’s Open Image Theorem [34], the image of the adelic Galois representation associated to $E_{0}/\mathbb{Q}$ has finite index in $\operatorname{GL}_{2}(\widehat{\mathbb{Z}})$. For $d$ fixed, it follows that there is a bound on the order of a cyclic subgroup of $E_{0}$ which can become $F$-rational over any number field $F$ of degree at most $d$. Thus there are only finitely many choices for $C$ such that $E_{\text{min}}\cong E_{0}/C$ over $\overline{\mathbb{Q}}$. ∎
@@
The minimal torsion curve within a fixed geometric isogeny class may or may not be unique (up to isomorphism over $\overline{\mathbb{Q}}$). For example, let $E_{0}/\mathbb{Q}$ be the elliptic curve with LMFBD label 38.b2 and let $\mathcal{E}$ denote its geometric isogeny class. Then Proposition 3.2 implies any minimal torsion curve for $X_{1}(5)$ must have $j$-invariant in $\mathbb{Q}$. By Lemma 3.1, the only other elliptic curve $E^{\prime}\in\mathcal{E}$ with $j$-invariant in $\mathbb{Q}$ has $j(E^{\prime})=-\frac{37966934881}{4952198}$. However, this curve does not give points on $X_{1}(5)$ of minimal degree. Thus $E_{0}$ is a unique minimal torsion curve for $X_{1}(5)$, up to $\overline{\mathbb{Q}}$-isomorphism.
On the other hand, let $E_{0}/\mathbb{Q}$ be the elliptic curve with LMFDB label 50.b1 and let $\mathcal{E}$ denote its geometric isogeny class. Then a similar argument shows that any elliptic curve $\mathbb{Q}$-isogenous to $E_{0}$ is a minimal torsion curve for $X_{1}(3)$, giving 4 distinct minimal torsion curves for this class.
4.2. Minimal Torsion Curves for Varying Modular Curves
If we allow $N$ to vary, there may be infinitely many elliptic curves (up to isomorphism over $\overline{\mathbb{Q}}$) within a fixed geometric isogeny class which are minimal for $X_{1}(N)$. For example, suppose the $\ell$-adic Galois representation of a non-CM elliptic curve $E/\mathbb{Q}$ is surjective, and let $C$ be a cyclic subgroup of $E$ of order $\ell^{k}$ for $k\in\mathbb{Z}^{+}$. The elliptic curve $E/C$ can be defined over the extension $\mathbb{Q}(C)$ of degree $\ell^{k-1}(\ell+1)$ and possesses a $\mathbb{Q}(C)$-rational cyclic $\ell^{k}$-isogeny. By Proposition 2.3, the curve $E/C$ gives a closed point on $X_{1}(\ell^{k})$ of degree $\ell^{2k-2}(\ell^{2}-1)/2.$ This is minimal for the geometric isogeny class of $E$ by Proposition 3.2. However, if $\mathcal{E}$ is a rational $\overline{\mathbb{Q}}$-isogeny class of non-CM elliptic curves, there always exists a minimal torsion curve for $X_{1}(\ell^{k})$ whose isogeny distance from a rational elliptic curve is bounded.
Proposition 4.3.
Let $\mathcal{E}$ be a rational $\overline{\mathbb{Q}}$-isogeny class of non-CM elliptic curves, and let $\ell$ be prime. There exists a constant $C=C(\mathcal{E},\ell)$ such for any $k\in\mathbb{Z}^{+}$ a point of least degree on $X_{1}(\ell^{k})$ coming from $\mathcal{E}$ can be associated to $j_{\text{min}}\in\mathcal{E}$ which is $d$-isogenous to a rational $j$-invariant for some $d\leq C$.
@@ 4.2. Minimal Torsion Curves for Varying Modular Curves
We note $\operatorname{im}\rho_{E_{0},\ell^{\infty}}$ has level $\ell^{k_{0}}$ for some $k_{0}\in\mathbb{Z}^{\geq 0}$ by Serre’s Open Image Theorem [34]. Replacing $k_{0}$ with a larger integer if necessary, we may assume $2k_{0}-2-d\geq 0$ if $\ell$ is odd and $2k_{0}-3-d\geq 0$ if $\ell=2$. Now, let $k$ be an integer with $k\geq k_{0}$. Then by Proposition 3.2, there exists $\alpha_{0}\in\mathbb{Z}^{+}$ such that the least degree of a closed point on $X_{1}(\ell^{k})$ associated to $E_{0}\in\mathcal{E}$ is
@@
Suppose first that $E_{0}$ is a minimal torsion curve for $X_{1}(\ell^{k})$ for sufficiently large $k$. Then by Proposition 4.1 there exists a finite collection of elliptic curves $E_{0},E_{1},\dots,E_{s}$ such that for any positive integer $n$ a point of least degree on $X_{1}(\ell^{n})$ coming from $\mathcal{E}$ can be associated to $E_{i}$ for some $0\leq i\leq s$. We may assume the isogeny $\varphi:E_{i}\rightarrow E_{0}$ is cyclic of degree $d_{i}$ by [14, Lemma A.1]. The result follows in this case with $C=\max\{d_{0},d_{1},\cdots,d_{s}\}$.
@@
From Proposition 4.3 we can immediately deduce the following corollary.
@@ 5. Results for non-CM classes and primes $\ell\geq 5$
The main result of this section is the following, which improves upon the divisibility conditions of [8, Proposition 4.1] for primes $\ell\geq 5$ and also characterizes the corresponding minimal torsion curves for points of odd degree. It proves Theorem 1.1 for $\ell\geq 5$.
@@
[…]
@@
Moreover, there exists a point of degree $\delta$ on $X_{1}(\ell^{k})$ associated to $j_{\text{min}}\in\mathcal{E}$ which is at most $\ell$-isogenous to a rational $j$-invariant. One can take $j_{\text{min}}\in\mathbb{Q}$ unless $\ell=7$ and $\mathcal{E}$ contains the elliptic curve with $j$-invariant $3^{3}\cdot 5\cdot 7^{5}/2^{7}$.
@@ 5.1. A Preliminary Result
We begin with a preliminary result concerning geometric isogeny classes containing an elliptic curve over $\mathbb{Q}$ with a rational cyclic isogeny. In this case, the result largely follows from Proposition 3.2 and prior work of Greenberg [25] and Greenberg, Rubin, Silverberg, and Stoll [24].
@@
$\delta\coloneqq\begin{cases}\deg(x)\cdot\ell^{2k-2}\text{ if $\ell\neq 5$ or $\mathcal{E}$ does not contain $E^{\prime}/\mathbb{Q}$ with a rational cyclic 25-isogeny},\\ \deg(x)\cdot 5^{\max(0,2k-3)}\text{ if $\ell=5$ and $\mathcal{E}$ contains $E^{\prime}/\mathbb{Q}$ with a rational cyclic 25-isogeny}.\end{cases}$
Moreover, there exists a point of degree $\delta$ on $X_{1}(\ell^{k})$ associated to $j_{\text{min}}\in\mathcal{E}$ with $j_{\text{min}}\in\mathbb{Q}$.
@@
divides the degree of any point on $X_{1}(\ell^{k})$ associated to $\mathcal{E}$, where $d=\operatorname{ord}_{\ell}([\operatorname{GL}_{2}(\mathbb{Z}_{\ell}):\operatorname{im}\rho_{E_{0},\ell^{\infty}}])$. By Greenberg [25, Theorems 1 and 2, Remark 4.2.1] and Greenberg, Rubin, Silverberg, Stoll [24], we have $d=0$. Since $\deg(X_{1}(\ell^{k})\rightarrow X_{1}(\ell))=\ell^{2k-2}$, lifts of $x$ on $X_{1}(\ell^{k})$ show this divisibility condition is best-possible with $j_{\text{min}}=j(E_{0})$.
Now, suppose $\ell=5$ and $\mathcal{E}$ contains $E^{\prime}/\mathbb{Q}$ with a rational cyclic 25-isogeny. Then by work of Greenberg [25, Theorem 2], we have $\operatorname{ord}_{5}([\operatorname{GL}_{2}(\mathbb{Z}_{5}):\operatorname{im}\rho_{E^{\prime},5^{\infty}}])=1$. By replacing $E^{\prime}$ with a curve $\mathbb{Q}$-isogenous, we may assume $E^{\prime}$ has two independent 5-isogenies. The index of the image of the 5-adic Galois representation is unchanged by [25, Proposition 2.1.1]. Then $\operatorname{im}\rho_{E^{\prime},5}$ is one of the following, and we consider each case separately:
[…]
  • 5Cs: By Proposition 3.2 and Remark 3.4, we see $2\cdot 5^{2k-3}$ divides the degree of any point on $X_{1}(5^{k})$ associated to $\mathcal{E}$. Note $E^{\prime}$ has two independent 5-isogenies with kernels $C_{1}$ and $C_{2}$. Then $E_{2}\coloneqq E^{\prime}/C_{2}$ has a rational cyclic 25-isogeny. By Proposition 2.3, the curve $E_{2}$ gives a closed point on $X_{1}(5)$ in degree dividing 2 and a closed point on $X_{1}(25)$ in degree dividing 10. Lifts of this point show the divisibility condition is best-possible for all $k$ with $j_{\text{min}}=j(E_{2})$. ∎

### v2

@@ 3.3. Image of Galois Representations Under Isogeny
    If $\operatorname{im}\rho_{E_{1},\ell^{\infty}}=\pi^{-1}(\operatorname{im}\rho_{E_{1},\ell^{k_{1}}})$ for some $k_{1}\in\mathbb{Z}^{+}$, then $\operatorname{im}\rho_{E_{2},\ell^{\infty}}=\pi^{-1}(\operatorname{im}\rho_{E_{2},\ell^{k_{1}+r}})$. Here, $\pi$ denotes the reduction map from $\operatorname{GL}_{2}(\mathbb{Z}_{\ell})$ to $\operatorname{GL}_{2}(\mathbb{Z}/\ell^{k_{1}}\mathbb{Z})$ or $\operatorname{GL}_{2}(\mathbb{Z}/\ell^{k_{1}+r\mathbb{Z}})$, respectively.
@@
If we can show that $M\in\operatorname{im}\rho_{E_{2},\ell^{k_{1}+r+1}}$, then $\operatorname{im}\rho_{E_{2},\ell^{\infty}}=\pi^{-1}(\operatorname{im}\rho_{E_{2},\ell^{k_{1}+r}})$ by, e.g., [BELOV, Proposition 3.5]. By the first paragraph, we may assume
@@
Suppose there exists an isogeny $\varphi:E\rightarrow E^{\prime}$ defined over $\overline{\mathbb{Q}}$ which is cyclic of order $\ell^{r}$. Thus there exists a point $P\in E$ of order $\ell^{r}$ such that $E^{\prime}\cong E/\langle P\rangle$. Let $F\coloneqq\mathbb{Q}(\langle P\rangle)$, and let $\psi:E\rightarrow E/\langle P\rangle$ be the induced $F$-rational isogeny. By Proposition 3.5 the $\ell$-adic Galois representation of $E/\langle P\rangle$, and thus the degrees of all closed points on $X_{1}(\ell^{k})$ associated to $E/\langle P\rangle$, is determined by $\operatorname{im}\rho_{E/F,\ell^{\infty}}$. Since $\operatorname{im}\rho_{E/F,\ell^{\infty}}$ can be obtained from $\operatorname{im}\rho_{E/\mathbb{Q},\ell^{\infty}}$ and the definition of $F$, the result follows. ∎
@@ 4. Properties of Minimal Torsion Curves
Let $\mathcal{E}$ be a rational $\overline{\mathbb{Q}}$-isogeny class of elliptic curves. In this section, we show that for a fixed positive integer $N$, there are finitely many minimal torsion curves in $\mathcal{E}$ for $X_{1}(N)$. In some cases, there is a unique minimal torsion curve, but this should not be expected in general (see Remark 4.2). Next, we suppose $\mathcal{E}$ is non-CM, [^2] Recall $\text{End}(E)\otimes\mathbb{Q}$ is an isogeny invariant, so the entire class $\mathcal{E}$ either is CM or non-CM. and let $E_{0}\in\mathcal{E}$ be an elliptic curve with rational $j$-invariant. Fix a prime $\ell$. In Theorem 4.3 we show that for any $k\in\mathbb{Z}^{+}$, there exists a minimal torsion curve $E\in\mathcal{E}$ for $X_{1}(\ell^{k})$ such that the degree of the isogeny from $E$ to $E_{0}$ is at most $C$, where $C$ is a constant that does not depend on $k$. This implies Theorem 4.3, as stated in the introduction.
@@ 4.1. Minimal Torsion Curves for Fixed Modular Curve
Let $d$ be the minimal degree of a point on $X_{1}(N)$ associated to $\mathcal{E}$, and let $j_{min}$ be the $j$-invariant of a minimal torsion curve for $\mathcal{E}$. Then $\mathbb{Q}(j_{min}):\mathbb{Q}]\leq d.$ If $\mathcal{E}$ is CM, then the conclusion follows as there are only finitely many CM $j$-invariants in an extension of bounded degree (since there are only finitely many imaginary quadratic fields—and hence imaginary quadratic orders—of a given class number [heilbronn], this is a consequence of [cox, Theorem 11.1 & Proposition 13.2]). So suppose $\mathcal{E}$ is non-CM.
Let $E_{min}\in\mathcal{E}$ be an elliptic curve with $j(E_{min})=j_{min}$. Since $\mathcal{E}$ is rational, there exists an elliptic curve $E_{0}/\mathbb{Q}\in\mathcal{E}$ with an isogeny $\varphi:E_{0}\rightarrow E_{min}$ defined over $\overline{\mathbb{Q}}$. We may assume $\varphi$ is cyclic of degree $n$ by [CremonaNajmanQCurve, Lemma A.1]. By [ClarkVolcanoes, Proposition 3.2] or [CremonaNajmanQCurve, Corollary A.5], the field of moduli of this isogeny is
$\mathbb{Q}(j_{min},j(E_{0}))=\mathbb{Q}(j_{min}).$
Thus $\varphi$ can be defined over $\mathbb{Q}(j_{min})$ by replacing $E_{0}$ and $E_{min}$ with appropriate twists, and in particular a twist of $E_{0}$ defined over this field has a rational cyclic $n$-isogeny. Since isogenies are twist invariant (i.e., if $E_{1},E_{2}$ are two elliptic curves over $F$ with the same $j$-invariant, then $E_{1}$ has an $F$-rational cyclic $n$-isogeny if and only if $E_{2}$ does), it follows that $E_{0}$ attains a rational cyclic $n$-isogeny over $\mathbb{Q}(j_{min})$, a number field of degree at most $d$.
[…]
@@
The minimal torsion curve within a fixed geometric isogeny class may or may not be unique (up to isomorphism over $\overline{\mathbb{Q}}$). For example, let $E_{0}/\mathbb{Q}$ be the elliptic curve with LMFDB label 38.b2 and let $\mathcal{E}$ denote its geometric isogeny class. Since there exists $P_{0}\in E_{0}(\mathbb{Q})$ of order 5, the least degree of a point on $X_{1}(5)$ associated to $\mathcal{E}$ is 1, and so any minimal torsion curve for $X_{1}(5)$ must have $j$-invariant in $\mathbb{Q}$. By Lemma 3.1, the only other elliptic curve $E^{\prime}\in\mathcal{E}$ with $j$-invariant in $\mathbb{Q}$ has $j(E^{\prime})=-\frac{37966934881}{4952198}$. However, this curve does not give points on $X_{1}(5)$ of minimal degree. Thus $E_{0}$ is a unique minimal torsion curve for $X_{1}(5)$, up to $\overline{\mathbb{Q}}$-isomorphism.
On the other hand, let $E_{0}/\mathbb{Q}$ be the elliptic curve with LMFDB label 50.b1 and let $\mathcal{E}$ denote its geometric isogeny class. Then a similar argument shows that any elliptic curve $\mathbb{Q}$-isogenous to $E_{0}$ is a minimal torsion curve for $X_{1}(3)$, giving 4 distinct minimal torsion curves for this class up to $\overline{\mathbb{Q}}$-isomorphism. Representatives for these curves can be found in the LMFDB isogeny class 50.b.
4.2. Minimal Torsion Curves of Varying Level
If we allow $N$ to vary, there may be infinitely many elliptic curves (up to isomorphism over $\overline{\mathbb{Q}}$) within a fixed geometric isogeny class which are minimal for $X_{1}(N)$. For example, suppose the $\ell$-adic Galois representation of a non-CM elliptic curve $E/\mathbb{Q}$ is surjective, and let $C$ be a cyclic subgroup of $E$ of order $\ell^{k}$ for $k\in\mathbb{Z}^{+}$. The elliptic curve $E/C$ can be defined over the extension $\mathbb{Q}(C)$ of degree $\ell^{k-1}(\ell+1)$ and possesses a $\mathbb{Q}(C)$-rational cyclic $\ell^{k}$-isogeny. By Proposition 2.3, the curve $E/C$ gives a closed point on $X_{1}(\ell^{k})$ of degree $\ell^{2k-2}(\ell^{2}-1)/2.$ This is minimal for the geometric isogeny class of $E$ by Proposition 3.2. However, if $\mathcal{E}$ is a rational $\overline{\mathbb{Q}}$-isogeny class of non-CM elliptic curves, there always exists a minimal torsion curve for $X_{1}(\ell^{k})$ whose isogeny degree to a rational elliptic curve is bounded.
Theorem 4.3.
Let $\mathcal{E}$ be a rational $\overline{\mathbb{Q}}$-isogeny class of non-CM elliptic curves, and let $\ell$ be prime. There exists a constant $C=C(\mathcal{E},\ell)$ such for any $k\in\mathbb{Z}^{+}$ a point of least degree on $X_{1}(\ell^{k})$ coming from $\mathcal{E}$ can be associated to $j_{min}\in\mathcal{E}$ which is $d$-isogenous to a rational $j$-invariant for some $d\leq C$.
@@ 4.2. Minimal Torsion Curves of Varying Level
We note $\operatorname{im}\rho_{E_{0},\ell^{\infty}}$ has level $\ell^{k_{0}}$ for some $k_{0}\in\mathbb{Z}^{\geq 0}$ by Serre’s Open Image Theorem [serre72]. Replacing $k_{0}$ with a larger integer if necessary, we may assume $2k_{0}-2-d\geq 0$ if $\ell$ is odd and $2k_{0}-3-d\geq 0$ if $\ell=2$. Now, let $k$ be an integer with $k\geq k_{0}$. Then by Proposition 3.2, there exists $\alpha_{0}\in\mathbb{Z}^{+}$ such that the least degree of a closed point on $X_{1}(\ell^{k})$ associated to $E_{0}\in\mathcal{E}$ is
@@
Suppose first that $E_{0}$ is a minimal torsion curve for $X_{1}(\ell^{k})$ for sufficiently large $k$. Then by Proposition 4.1 there exists a finite collection of elliptic curves $E_{0},E_{1},\dots,E_{s}$ such that for any positive integer $n$ a point of least degree on $X_{1}(\ell^{n})$ coming from $\mathcal{E}$ can be associated to $E_{i}$ for some $0\leq i\leq s$. We may assume the isogeny $\varphi:E_{i}\rightarrow E_{0}$ is cyclic of degree $d_{i}$ by [CremonaNajmanQCurve, Lemma A.1]. The result follows in this case with $C=\max\{d_{0},d_{1},\cdots,d_{s}\}$.
@@
From Theorem 4.3 we can immediately deduce the following corollary.
@@ 5. Results for non-CM classes and primes $\ell\geq 5$
The main result of this section is the following, which proves Theorem 1.2 for $\ell\geq 5$. It refines [OddDegQCurve, Proposition 4.1], which gives divisibility conditions which hold across all rational non-CM classes $\mathcal{E}$ and does not address whether they are best-possible.
@@
  • If $\ell=7$ and $\mathcal{E}$ contains $E^{\prime}/\mathbb{Q}$ with $\operatorname{im}\rho_{E^{\prime},7}\in\{7B.1.1,7B.1.6,7B.6.1\}$, then $\delta=7^{2k-2}$.
@@
Moreover, there exists a point of degree $\delta$ on $X_{1}(\ell^{k})$ associated to $j_{min}\in\mathcal{E}$ which is at most $\ell$-isogenous to a rational $j$-invariant. One can take $j_{min}\in\mathbb{Q}$ unless $\ell=7$ and $\mathcal{E}$ contains the elliptic curve with $j$-invariant $3^{3}\cdot 5\cdot 7^{5}/2^{7}$.
@@ 5.1. A Preliminary Result
We begin with a preliminary result concerning geometric isogeny classes $\mathcal{E}$ containing an elliptic curve over $\mathbb{Q}$ with a rational cyclic isogeny. When this occurs, we will say $\mathcal{E}$ gives a point in $X_{0}(\ell)(\mathbb{Q})$. In this case, the result is largely a consequence of the following theorem and Proposition 3.2.
Theorem 5.2 (Greenberg [greenberg2012], Greenberg, Rubin, Silverberg, and Stoll [greenberg2014]).
Let $E/\mathbb{Q}$ be a non-CM elliptic curve with a $\mathbb{Q}$-rational cyclic isogeny of prime degree $\ell$. By choice of basis, we may assume $\operatorname{im}\rho_{E,\ell^{\infty}}$ is a subgroup of $\operatorname{GL}_{2}(\mathbb{Z}_{\ell})$.
  1. (1)
    If $\ell\geq 7$, then $\operatorname{im}\rho_{E,\ell^{\infty}}$ contains a Sylow pro-$\ell$ subgroup of $\operatorname{GL}_{2}(\mathbb{Z}_{\ell})$.
  2. (2)
    Let $\ell=5$. If none of the elliptic curves in the $\mathbb{Q}$-isogeny class of $E$ has two independent isogenies of degree 5, then $\operatorname{im}\rho_{E,5^{\infty}}$ contains a Sylow pro-$5$ subgroup of $\operatorname{GL}_{2}(\mathbb{Z}_{5})$. Otherwise the index of $\operatorname{im}\rho_{E,5^{\infty}}$ in $\operatorname{GL}_{2}(\mathbb{Z}_{5})$ is divisible by 5, but not by 25.
Proof.
If $\ell\geq 11$ or $\ell=5$, this follows from [greenberg2012, Theorems 1 & 2]. Note if $\ell\geq 11$ or if $\ell=5$, the assumption in [greenberg2012, Theorems 1] holds, as explained in [greenberg2012, Remark 4.2.1, p. 1186–1187]. For $\ell=7$, this follows from [greenberg2014, Theorem 5.5]. ∎
@@
[…]
Moreover, there exists a point of degree $\delta$ on $X_{1}(\ell^{k})$ associated to $j_{min}\in\mathcal{E}$ with $j_{min}\in\mathbb{Q}$.
@@
divides the degree of any point on $X_{1}(\ell^{k})$ associated to $\mathcal{E}$, where $d=\operatorname{ord}_{\ell}([\operatorname{GL}_{2}(\mathbb{Z}_{\ell}):\operatorname{im}\rho_{E_{0},\ell^{\infty}}])$. By Theorem 5.2, we have $d=0$. Since $\deg(X_{1}(\ell^{k})\rightarrow X_{1}(\ell))=\ell^{2k-2}$, lifts of $x$ on $X_{1}(\ell^{k})$ show this divisibility condition is best-possible with $j_{min}=j(E_{0})$.
Now, suppose $\ell=5$ and $\mathcal{E}$ contains $E^{\prime}/\mathbb{Q}$ with a rational cyclic 25-isogeny. Then by Theorem 5.2, we have $\operatorname{ord}_{5}([\operatorname{GL}_{2}(\mathbb{Z}_{5}):\operatorname{im}\rho_{E^{\prime},5^{\infty}}])=1$. By replacing $E^{\prime}$ with a curve $\mathbb{Q}$-isogenous, we may assume $E^{\prime}$ has two independent 5-isogenies. The index of the image of the 5-adic Galois representation is unchanged by [greenberg2012, Proposition 2.1.1]. Then $\operatorname{im}\rho_{E^{\prime},5}$ is one of the following, and we consider each case separately:
  • 5Cs.1.1, 5Cs.1.3, or 5Cs.4.1: Replacing $E^{\prime}$ with a quadratic twist if necessary, we may assume $\operatorname{im}\rho_{E^{\prime},5}=5Cs.1.1$. By Proposition 3.2 and Remark 3.4, we see $5^{\max(0,2k-3)}$ divides the degree of any point on $X_{1}(5^{k})$ associated to $\mathcal{E}$. The curve $E^{\prime}$ has a subgroup $C_{1}$ generated by a rational point $P$ of order 5 and an independent rational cyclic subgroup $C_{2}$ of order 5. Then $E_{2}\coloneqq E^{\prime}/C_{2}$ has a rational cyclic 25-isogeny, and the image of $P$ is a point of order 5 in $E_{2}(\mathbb{Q})$ lying in its kernel. Thus the image of the isogeny character $\chi:\operatorname{Gal}_{\mathbb{Q}}\rightarrow(\mathbb{Z}/25\mathbb{Z})^{\times}$ lands in the subgroup $\{a:a\equiv 1\pmod{5}\}$ and so has order dividing 5. It follows that $E_{2}$ attains a rational point of order 25 in an extension of degree dividing 5. By Proposition 2.6, there are lifts of this point on $X_{1}(5^{k})$ of degree at most $5^{2k-3}$ for all $k\geq 2$. Thus the condition is best-possible for all $k$ with $j_{min}=j(E_{2})$.
  • 5Cs: By Proposition 3.2 and Remark 3.4, we see $2\cdot 5^{2k-3}$ divides the degree of any point on $X_{1}(5^{k})$ associated to $\mathcal{E}$. Note $E^{\prime}$ has two independent 5-isogenies with kernels $C_{1}$ and $C_{2}$. Then $E_{2}\coloneqq E^{\prime}/C_{2}$ has a rational cyclic 25-isogeny. By Proposition 2.3, the curve $E_{2}$ gives a closed point on $X_{1}(5)$ in degree dividing 2 and a closed point on $X_{1}(25)$ in degree dividing 10. Again by considering lifts of this point with degree bounds given by Proposition 2.6, we see the divisibility condition is best-possible for all $k$ with $j_{min}=j(E_{2})$. ∎